OpenStudy (anonymous):

₃⁄₄(2x+y)=9 ⅝(2x=11y)=20 Solve question 6.2 above graphically ,i.e, plotting the lines on the same set of axes and determining the point of intersection. Does your answer agree with that obtained above? can someone help me plz

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

simultaneous equations

OpenStudy (anonymous):

expand brackets

OpenStudy (anonymous):

1.5x+0.75y=9 1.25x+6.875y=20

OpenStudy (anonymous):

y+2x=12 --> y=-2x+12 y=2/11x+32/11

OpenStudy (anonymous):

point of intersection, merge y's together

OpenStudy (anonymous):

-2x+12=2/11x+32/11

OpenStudy (anonymous):

24/11x=100/11

OpenStudy (anonymous):

24x=100 --> x=25/6

OpenStudy (anonymous):

that's your x co-ord

OpenStudy (anonymous):

plug into any of the two equations to find y, and that's your co-ord

OpenStudy (anonymous):

jeeezzzz u r so good at it!!!

OpenStudy (anonymous):

medal please :]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

can i replace x co-ord in this equation to find y?25x+6.875y=20